4/3 x -2 1/2= -6 1/4 Solve
Answer by lwsshak3(11628)

4/3 x -2 1/2= -6 1/4 Solve
change to improper fractions
4x/3-5/2=-25/4
LCD: 12
multiply each term by LCD
16x-30=-75
16x=-45
x=-45/16